MAT 125, CALCULUS A, SUNY AT STONY BROOK, Spring 2005

Text: Single variable calculus: concepts and contexts, James Steward. Syllabus: We will cover approximately Chapters 1,2,3,4. A more detailed syllabus will

be provided. This includes some preliminary material, limits, differentiation and applications.

Classdays. Classes begin on Monday, January 24 and end on Monday, May 9. Holidays: There are no classes during spring break, March 21-26. There are also no

classes during Monday, April 25, for the observance of Passover.

Lecture times and locations:

day  time  room  instructor
LEC 1 MWF 10:40pm- 11:35am ENGINEERING 143  Reiris, Martin
R01  M  8:30am-9:25am  PHYSICS P116  Branson, Mark
R02  Th  3:50pm-4:45pm  CHEMISTRY 123  Holmes, Kristin
R03  Th  8:20am- 9:15am  OLD CHEM 138  van Coevering, Craig
R04  Tu  2:20pm-3:15pm  S B UNION 226  Liu, Yuan
R05  Th  11:20am-12:15am OLD CHEM 138  Shu, Yu-Jen
LEC 2 TuTh 12:50pm-2:10pm ENGINEERING 145  Bishop, Christopher
R06  w  11:45am-12:40pm PHYSICS P112  Chen, He-Wei
R07  F  11:45am-12:40pm PHYSICS P112  Thind, Jaimal
R08  W  9:35am-10:30pm SOCBEHAV SCI S218  Roberts, Amy
R09  W  8:30am-9:25am  OLD CHEM 138  Branson, Mark
R10  F  8:30am-9:25am  PHYSICS P116  Cabrera-Ocanas, Carlos
LEC 3 MW 2:20pm- 3:40pm ENGINEERING 145  Kumpel, Paul
R11  M  9:35am-10:30am  PHYSICS P118  Roberts, Amy
R12  F  10:40am-11:35am PYSICS P118  Chen, Je-Wei
R13  Tu  2:20pm-3:15pm  LGT ENGR LAB 152  Nam, Young Woo
R14  Tu  11:20am-12:15pm OLD CHEM 138  Chance, Michael
R15  Th  11:20amm-12:15pm PHYSICS P117  Chance, Michael
ELC 4 TuTh 6:50pm- 8:10pm HARRIMAN HALL 116 Suvaina,Ioana

There is no recitation for ELC 4. The grader for ELC 4 is Nam, Young Woo.

Instructor  Office Office hours  email
Bishop  4-112 TuTh 11:30-12:30am  bishop@math.sunysb.edu

Kumpel  3-120 MW 4-5pm  pkumpel@math.sunysb.edu

Suvaina  3-116 Tu 5:30-6:30 in 3-116, Th 5:00-6:00 in MLC sio@math.sunysb.edu
Reiris  2-112 W 11:45-12:45 in 2-112, F 12:00-1:00 in MLC martinr@math.sunysb.edu

Roberts  3-119  aroberts@math.sunysb.edu

Nam  2-117 MW 2:30-3:30 in 2-117, M 4-5 in MLC  ynam@math.sunysb.edu

Chance  mchance@math.sunysb.edu
Holmes  M 2:30-4:30 in MLC  kristinholmes@optonline.net
vanCoevering  craig@math.sunysb.edu

Liu  4-118 Tu 3:30-5:30 in MLC, 12:30-1:30 in 4-118  yliu@math.sunysb.edu

Thind  jthind@math.sunysb.edu
Shu  W 1:30-3:30pm in MLC  yjshu@math.sunysb.edu
Branson  mbranson@math.sunysb.edu

Chen  2-106 W 1-3 in 2-106; Th 1-2 in MLC  jwchen@math.sunysb.edu

Cabrera-Ocanas  carlos@math.sunysb.edu

Homework. A list of problems from the text will be handed out. The recitation instructor will assign, collect and return homework. In general, problems are due the week after the section has been covered in lecture.

Examinations. There will be two midterm tests: Monday, Febuary 21 and Tuesday, March 29???? (as of this writing the date of the second exam still needs to be confirmed). These will be given at night (8:30pm-10:00pm) in a location which will be annouced in class. Please make sure you have no conflicts with these exams times. The final exam is on Monday, May 16 at 11:00-1:30. The location will be announced in class during the last week of the semester. No calculators, notes or books will be allowed during exams. Bring your student ID to all exams.

Grades. Grades are computed using: Recitation 10%, Midterms 25% each, Final 40%.

Webpage: Course information is posted at:

http://www.math.sunysb.edu/~bishop/classes/math125.S05/math125.html

Disabilities. If you have any condition, such as physical or mental disability, that will make it difficult to complete the course or that will require extra time on examinations, notify your lecturer during the first two weeks of class so that appropriate arrangements can be made.

Math Learning Center: This is in a new location: S-240A in the basement of the math building. MLC provides help with pre-calculus and calculus material. It is staffed most days and some evenings; an exact schedule will be posted outside the room.

Drop dates: Last day for Add/Drop is January 30. Last day to switch to lower numbered math course is March 18. The last day to withdraw from a course is April 1.